find the equation of line which is perpendicular to line y=5x+9 and has y-intercept=6
Answer by gc2000(22)

summary:
slope = m and y-intercept = b when line is expressed as y=mx+b
Parallel lines have the same slope, perpendicular lines have slopes that are "negative reciprocals" (for example, if slope of a line is 3, then -1/3 is the slope for all perpendicular lines; if slope is 1/2, perpendicular slopes are -2, etc.).


y=5x+9
since slope is 5, all perpendicular lines have slope = -1/5
since y-intercept is 6, "b"=6 in formula: y=mx+b


substitute in -1/5 and 6 into y=mx+b:
y=(-1/5)x+6